OpenStudy (anonymous):

\[y=\frac 1 x\] If the region R{(x,y)|x>=1,0<=y<= (1/x)} is rotated about the x-axis, the volume of the resulting solid is infinite. Show that the surface area is infinite. (aka Gabriel's horn)

OpenStudy (anonymous):

\[\int_{1}^{\infty} 2\pi\frac 1 x \sqrt(1+(\frac{-1}{x^2})^2)dx\]

OpenStudy (anonymous):

alright continue

OpenStudy (blockcolder):

Lemme clean that up. \[2\pi \int_1^\infty \frac{1}{x}\sqrt{1+\frac{1}{x^4}}\ dx\]

OpenStudy (anonymous):

alright kep going

OpenStudy (anonymous):

hint do someting with the radical

OpenStudy (anonymous):

substitution?

OpenStudy (anonymous):

not quite , create a common denominator

OpenStudy (anonymous):

\[2\pi\int_{1}^{\infty} \frac 1 x \sqrt(1+\frac{1}{x^4})dx\] \[\sqrt(1+\frac{1}{x^4})\] \[\sqrt(\frac{1+x^4}{x^4})\] \[ \frac{\sqrt(1+x^4)}{x^2}\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

continue

OpenStudy (anonymous):

\[2\pi\int_{1}^{\infty} \frac 1 x \frac{\sqrt(1+x^4)}{x^2}dx\] \[2\pi\int_{1}^{\infty} \frac{\sqrt(1+x^4)}{x^3}dx\]

OpenStudy (anonymous):

OH MY GOD! \[2\pi\int_{1}^{\infty} \frac{\sqrt(1+x^4)}{x^3}dx\] u=1+x^4 du=4x^3 (1/4)du=x^3

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I realized that I can finally make the substitution to solve the integral

OpenStudy (anonymous):

oh hah

OpenStudy (anonymous):

that's not right though because the x^3 is in the denominator

OpenStudy (anonymous):

\[\frac\pi 2\int_{1}^{\infty} \frac{\sqrt(u)}{x^3}du\] what about the \[x^3\] that's in the denominator?

OpenStudy (anonymous):

trig sub =P

OpenStudy (anonymous):

do you know this?

OpenStudy (anonymous):

I thinks so ...lets see \[\]

OpenStudy (anonymous):

\[2\pi\int_{1}^{\infty} \frac{sec^2}{x^3}dx\] \[sec^2\] because 1+tan^4 =sec^4 x=tan^4 x dx=4 tan^3 x sec^2 x

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

\[2\pi\int_{1}^{\infty} \frac{sec^2\theta}{(tan^4 \theta)^3} 4tan^3 \theta sec^2\theta d\theta\]

OpenStudy (anonymous):

\[8\pi\int_{1}^{\infty} \frac{sec^4\theta}{tan^4 \theta} d\theta\]

OpenStudy (anonymous):

let x=atan(x) dx=asec^2x

OpenStudy (anonymous):

so what I have written is wrong?

OpenStudy (anonymous):

\[2\pi\int_{1}^{\infty} \frac{\sqrt(1+tanx^4)}{tanx^3}sec^2xdx\]

OpenStudy (anonymous):

i'm just looking to make sure x=tan(x) dx=sec^2x \[\sqrt{a^2+x^2}=\sqrt{1+tan^2x}=\sqrt{sec^2x}=sec(x)\]

OpenStudy (anonymous):

the numerator seems to be \[sec^4x\]

OpenStudy (anonymous):

is the denominator ok?

OpenStudy (anonymous):

wait it should be \[x^2=tan(x)\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

\[\pi\int_{1}^{\infty} \frac{\sqrt(1+tan^2 x)}{xtanx}sec^2xdx\] dx=.5sec^2 x?

OpenStudy (anonymous):

i'm not so sure now lol

OpenStudy (anonymous):

\[8\pi\int_{1}^{\infty} \frac{sec^4\theta}{tan^4 \theta} d\theta\] this look pretty...

OpenStudy (dumbcow):

\[x = \sqrt{\tan u}\] \[dx = \frac{\sec^{2} u}{2\sqrt{\tan u}}\] \[\rightarrow 2\pi \int\limits_{1}^{\infty} \frac{\sec u}{\tan u \sqrt{\tan u}}*\frac{\sec^{2} u}{2\sqrt{\tan u}} du\] \[= \pi \int\limits_{1}^{\infty}\frac{\sec^{3} u}{\tan^{2} u} du\]

OpenStudy (anonymous):

ahh so i was right -.-

OpenStudy (anonymous):

sorry sofia i'm just really busy atm

OpenStudy (anonymous):

lol no prob.

OpenStudy (anonymous):

thanks smartcow

OpenStudy (anonymous):